Sam has 15 identical rectangular windows in his home. The length of each window is twice the width, and the area of each window is 8 square feet. Sam wants to put weather stripping around the perimeters of all the windows.
If the weather stripping costs $0.25 per foot, how much will he spend?

Answer by Cromlix(4381)

Hi there,
15 windows.
Length = 2 x width.
Make width = x
Length = 2x
Area of window = width x length
8 = x(2x)
2x^2 = 8
x^2 = 4
x = √4
x = 2 feet.
Width = 2 feet
Length = 4 feet.
Perimeter = 2 x Width + 2 x Length
Perimeter = 2(2) + 2(4)
Perimeter of window = 12 feet
15 windows = 15 x 12
= 180 feet
180 x $0.25
= $45
Hope this helps :-)